Question 7. There are 40 students in the Statistics class; of them, 15 are girls and 25 are boys. The teacher wants to select randomly a group of ten for a volunteer activity. What is the probability of making the volunteer group with 5 girls and 5 boys? *
1
Expert's answer
2020-04-27T17:44:18-0400

Their order does not matter, so we can use formula:

P=(C(15,5)*C(25,5))/C(40,10)=(3003*53130)/847660528=0.1882